### v2

# Synth2: Boosting Visual-Language Models with Synthetic Captions and Image Embeddings

Abstract
The creation of high-quality human-labeled image-caption datasets presents a significant bottleneck in the development of Visual-Language Models (VLMs). In this work, we investigate an approach that leverages the strengths of Large Language Models (LLMs) and image generation models to create synthetic image-text pairs for efficient and effective VLM training. Our method employs a pretrained text-to-image model to synthesize image embeddings from captions generated by an LLM. Despite the text-to-image model and VLM initially being trained on the same data, our approach leverages the image generator’s ability to create novel compositions, resulting in synthetic image embeddings that expand beyond the limitations of the original dataset. Extensive experiments demonstrate that our VLM, finetuned on synthetic data achieves comparable performance to models trained solely on human-annotated data, while requiring significantly less data. Furthermore, we perform a set of analyses on captions which reveals that semantic diversity and balance are key aspects for better downstream performance. Finally, we show that synthesizing images in the image embedding space is 25% faster than in the pixel space. We believe our work not only addresses a significant challenge in VLM training but also opens up promising avenues for the development of self-improving multi-modal models.
keywords:
synthetic data, visual language models, multimodal self-improvement, open-ended learning1 Introduction
Visual-language models (VLMs) are quickly emerging as powerful tools for understanding visual and textual information. Their ability to combine these two modalities holds immense promise for applications ranging from image captioning to visual question answering. While VLMs hold significant potential, their performance is often constrained by limited data availability. Recent breakthroughs demonstrate that pre-training VLMs on larger image-text pair datasets leads to significant improvements in downstream tasks (Li et al., 2022b; Hu et al., 2022). However, creating such datasets poses several challenges such as scarcity of paired data, potentially noisy nature of data sourced from the internet (e.g., LAION (Schuhmann et al., 2021)), high curation costs, low diversity and high imbalance in semantics. These factors often lead to laborious filtering and extended training times due to low signal-to-noise ratio, thus increasing overall resource consumption (Figure 1A).
This work tackles these limitations by introducing an efficient approach that leverages pre-trained generative text and image models to create synthetic paired data for VLMs (see Figure 1B). Our approach uniquely synthesizes both text and images, overcoming the reliance on real-world data and addressing the challenges of scarcity, cost, and noise.
Additionally, the proposed framework operates at both the pixel and embedding levels (see Figure 1C), enabling us to train the VLM from either real images or synthetic image embeddings. This bypasses the need for pixel-space rendering for both the image generator and the VLM encoder. This paradigm shift significantly reduces memory overhead and resource consumption while maintaining the quality of synthetic training data.
While synthetic data generation has been explored for various computer vision tasks such as image segmentation, optical flow prediction, or image classification (Mishra et al., 2022; Greff et al., 2022; Azizi et al., 2023; Fan et al., 2023; Li et al., 2023c), its application to both visual and textual modalities within VLM training is a significant advancement. Furthermore, a potential pitfall in studying the impact of synthetic data in training new models lies in the inherent advantage of off-the-shelf image generators. These models, trained on massive image-text datasets, may inadvertently equip VLMs with knowledge already embedded in the image generators’ own pre-training data. This overlap could mask the true effectiveness of synthetic images, leading to improvements that might be equally achievable by simply training VLMs on the original image generator dataset.
To isolate the unique contribution of synthetic images, we propose a controlled study. Both the text-to-image generator and the VLM undergo pre-training on the exact same dataset. This alignment ensures that any subsequent improvements observed in the VLM after fine-tuning on synthetic data can be confidently attributed to the novel compositions and representations generated, rather than pre-existing knowledge inherited from the original dataset.
To sum up, our research explores the synergy between VLMs and text-to-image generation models, demonstrating a powerful framework for overcoming data limitations. This paper presents the following key contributions:
Fully Synthetic Data Creation: We introduce the first VLM training process that utilizes a fully synthetic dataset of high-quality text-image embedding pairs. These pairs are generated by pre-trained generative models, circumventing the need for massive amounts of real-world data for either modality.
Efficient Embedding Space Generation: Our method works with images in embedding space, avoiding costly image rendering and dramatically improving efficiency, without comprising performance.
Fair Evaluation Through Control: By pre-training a text-to-image model on the same dataset used for VLM training, instead of using a large, off-the-shelf model, we prevent knowledge transfer from a model trained on vast, external image-text pairs. This fosters a fair and focused assessment of the benefits of synthetic data in isolation.
Demonstrated Effectiveness: Experiments will showcase significant performance gains in 3 domains (scene description, scene understating QA and Extrernal Knowledge QA) when using our synthetic data, assessed against real and synthetic-data baselines.
In conclusion, we offer insights into the future of VLM training, highlighting the potential for creating customized datasets, open-ended training and self-improving multimodal models.
2 Related Work
Generator | Method | Generator Model | Generated Set | Caption Class | Caption Type | Evaluation Setting | |
Canonical Concept Mapping | Sharifzadeh et al. (2021) | Linear | Scene Graph Embedding | Scene Graphs | |||
Sharifzadeh et al. (2022) | Complex Text | Human Generated | SG Classification | ||||
Mishra et al. (2022) | Mixed pairs such as | ||||||
Greff et al. (2022) | (Segmentation, Images) | ||||||
Simulation/Rendering Engine | Zheng et al. (2020) | Mix | (Optimal Flow, Videos) | N/A | N/A | Mix | |
de Melo et al. (2022) | (Depth Maps, Images) | ||||||
Cascante-Bonilla et al. (2023) | Mix | (Captions, Images) | Complex Text | Rule-based | Vision Encoder | ||
Azizi et al. (2023) | SD | ||||||
Off-the-shelf Image Generator | Fan et al. (2023) | Imagen | Images | Single Word | ImageNet Classes | Classifier | |
MUSE | |||||||
Li et al. (2023c) | SD | Images | Complex Text | Human Generated | VLM | ||
Controlled Image Generator | Synth2 | MUSE | Text, Embeddings & Images | Complex Text | Human & LLM Generated | VLM |
Visual-language models (VLMs) pretrained on large-scale image-text data, primarily sourced from the web, have shown remarkable capabilities across numerous vision-language tasks. These tasks include image captioning, visual question answering, and few-shot learning (Alayrac et al., 2022; Chen et al., 2022; Li et al., 2022b). Architecturally, VLMs commonly employ an image encoder coupled with a large language model (LLM). Researchers explore diverse image encoder designs, ranging from convolutional neural networks (CNNs) (Alayrac et al., 2022) to transformer-based architectures (Chen et al., 2022). Additionally, the choice between pretraining the image encoder on separate image datasets (Alayrac et al., 2022) and training it from scratch alongside the VLM (Tsimpoukelli et al., 2021) remains an active area of investigation. Pretraining VLMs involves various objectives, including supervised learning approaches, as well as contrastive methods that focus either on aligning image and text representations (Radford et al., 2021) or on aligning different image representations (Chen et al., 2020b, a; Grill et al., 2020). Similar diversity exists in LLM choices, with both encoder-decoder (Chen et al., 2022) and decoder-only (Alayrac et al., 2022) architectures being utilized. The strategies for combining information extracted by the image encoder and the language model also exhibit substantial variation. Despite the success achieved with web-scraped datasets, there’s a growing emphasis on designing VLM architectures that facilitate training using synthetically generated data. This capability holds particular significance for applications where data scarcity or resources availability pose significant challenges.
Generation of synthetic data for training machine learning models remains a highly active area of research. While numerous studies (Mishra et al., 2022; Cascante-Bonilla et al., 2023; Greff et al., 2022; Zheng et al., 2020; de Melo et al., 2022) explore the use of model-based rendering engines or simulators, the remarkable advancements in high-quality image generators have ignited a surge of interest in leveraging generative models for creating training data. This trend has significantly impacted a wide range of computer vision tasks, including semantic segmentation (Li et al., 2022a; Ros et al., 2016; Baranchuk et al., 2021; Tritrong et al., 2021; Li et al., 2021; Chen et al., 2019), human motion understanding (Varol et al., 2017; Ma et al., 2022; Guo et al., 2022), and more recently image classification (Azizi et al., 2023; Fan et al., 2023). Our work investigates the potential of data-driven generative models within the domain of visual-language models (VLMs). We focus on their application in downstream tasks, where the ability to process complex scene is important. From this aspect, our approach is closest to a concurrent work by Li et al. (2023c) where the goal is to replace faulty images in a captioning pipeline with their synthetic versions. However, our work distinguishes itself from Li et al. (2023c) by achieving strong performance while utilizing 40 times less paired data and only a quarter of the parameters. This demonstrates the potential for achieving both performance and efficiency through our proposed approach. Furthermore, unlike the mentioned works (Azizi et al., 2023; Fan et al., 2023; Li et al., 2023c), we avoid using an off-the-self image generator trained on larger datasets. This approach prevented knowledge transfer into the VLM from a model trained on large amounts of paired data, which is critical as such knowledge transfer would have obscured our ability to scientifically assess the unique contribution of synthetic data.
Our work can also be likened to the concept of cycle consistency (Zhu et al., 2017) in visual-language models. This principle, where image-to-text and text-to-image conversion can be employed as a form of synthetic data training, albeit with extra gradients throughout the network, exploring image-text cycle consistency during sampling (Li et al., 2023a) or training (Li et al., 2023b) has been explored in recent works, demonstrating promising results.
Finally, we emphasize the efficiency of pipelines that seamlessly connect data generation and model training. While most prior work has concentrated on generating images in pixel space, we investigate the generation of image embeddings that can be directly integrated into the VLM. Our approach aligns with recent work in the scene graph classification domain (Sharifzadeh et al., 2021, 2022), which has explored synthetic scene graph embedding generation using pretrained canonical class representations.
Table 1 provides a taxonomy of some of the mentioned related work to clarify the differences better. As shown, our study is the first of its kind to scientifically explore the application of synthetic data generation within a VLM training pipeline and in particular using image embeddings and by generating synthetic captions.
3 Synth2
Given the synthetic generation of text and then images, we refer to our method as Synth2. Synth2 is a pipeline for training VLMs using generative models in addition to collected human data. In this section we introduce different components of this pipeline, namely Caption Generation, Image Generation and the full Synth2 model (see Figure 3 for details).
3.1 Synthetic Caption Generation
We leverage the generative capabilities of LLMs to create synthetic captions. To ensure a wide variety of captions, we adopt a class-based prompting approach. First, we randomly select a class from the ImageNet21k dataset (Ridnik et al., 2021). The LLM (Gemini Pro (Team et al., 2023)) is then prompted with the following text:
Make up a human-annotated description of an image that contains the following object: . The caption should be around 30-40 words long. Describe the different components of the scene in an objective and unbiased way. Do not add subjective judgments about the image, it should be as factual as possible. Do not use fluffy, poetic language. Respond only with the caption itself, beginning with “This is an image of”.
where we replace “[object]” with the randomly selected class. This class-based prompting encourages the LLM to generate captions that cover a broad range of visual concepts. Figure 2 shows some samples from the generated captions.
3.2 Image Generation
Our image generator architecture is similar to Chang et al. (2023) and was chosen for its superior inference efficiency due to parallel decoding and its use of discrete image tokens that dramatically reduces the number of sampling iterations compared to auto-regressive models. Additionally, MUSE leverages a transformer architecture, further enhancing efficiency (Chang et al., 2023). As discussed in Section 1, to mitigate the potential for knowledge overlap between off-the-shelf image generators and VLMs, we deviate from standard practice. Instead of employing pre-trained image generators, we intentionally pre-train both our text-to-image generator and VLM on an identical dataset. This controlled setup ensures that any performance gains observed in the VLM after fine-tuning on the synthetic data are directly attributable to the unique contributions of the generated images, rather than inherited knowledge from the original dataset. Hence, we pretrain an image generator on a image-caption dataset that will also be utilized for VLM pretraining (i.e. Conceptual Captions v2 (Changpinyo et al., 2021) see Section 4).
3.2.1 Training
To pretrain our image generator given paired image and texts, we embed the texts with a pre-trained language model and the images with a pre-trained VQ-GAN (Esser et al., 2021) (refer to A.1.1 in the appendix for the details). The pre-trained language model is the same used in our VLM and it is a reduced version of Chinchilla (Hoffmann et al., 2022) using 400 million parameters. Following the approach in Chang et al. (2023), we apply a masking procedure with a cosine scheduling function. This function gradually increases the masking ratio during training, biasing the model towards learning from images with larger masked regions. On average, around 64% of the VQ tokens are replaced with a special “dropped token”. The noisy VQ tokens and the embedded text are then fed into a transformer model. The model’s objective is to predict the original, unmasked VQ tokens. The loss function is the cross-entropy between the predicted VQ tokens and the masked ones:
where denotes the VQ tokens computed from the ground-truth image , denotes the ground-truth caption, and denotes the indices of the masked tokens. Our text-to-image generator was trained only on 10.1 millions text-image pairs from Conceptual Captions V2 (Changpinyo et al., 2021) (see Section 4 for details).
3.2.2 Inference
During inference, we use an iterative procedure to generate images. At first iteration, the predicted VQ tokens are all initialized with the “dropped token” representing the masked entries. At each upcoming iteration, the text embedding and the predicted VQ tokens until that stage are fed into the transformer and the model predicts the masked tokens. A cosine schedule determines which masked tokens have the highest prediction confidence during decoding. These tokens are unmasked, progressively reducing the masked token set. This process repeats for 24 iterations, gradually refining the generated VQ-tokens. Additional details for our text-to-image generator model and examples of generated images, given ground truth annotations compared to the original ones, are reported in A.1.3 in the appendix.
3.3 Synth2 VLM architecture
As described in Section 3.2, our text-to-image model generates image embeddings through an iterative caption-conditioned denoising process. A VQ-GAN decoder could be used to convert these tokens into human-viewable images. However, we emphasize that the image embeddings themselves provide a compressed representation of an image, making the decoder step unnecessary for certain applications.
Therefore, to enhance efficiency for training the VLM from synthetic data, we design the VLM to enable the bypass of pixel-space processing. This is achieved by setting its vision encoder to be identical to the VQ-GAN backbone used for our image generator. This enables seamless interaction between the VLM and synthetically generated image embeddings and eliminates the computationally expensive decoding stage of VQ-GAN for the image generator and the encoding stage for the VLM when training on synthetic data. At the same time, it still allows us to train from images in pixel space when using human annotated data. This design feature is particularly important since in order to curtail model collapse effects (Shumailov et al., 2023), we need to mix synthetic and real data during finetuning (Gerstgrasser et al., 2024).
Overall, our VQ-based design removes the need for costly conversions to and from pixel space, resulting in significant time and disk space savings as will be shown in Section 4. To harness the efficiency of discrete tokens during training and sampling, while maintaining the rich image understanding provided by soft embeddings, we convert each discrete token to its soft embedding using the codebook before feeding them to the VLM.
On top of the VQ backbone in our VLM, we incorporate a Perceiver Resampler component (Jaegle et al., 2021) to cross-attend to the VQ tokens extracted from our backbone. Similar to Flamingo Alayrac et al. (2022), we employ a frozen language model to cross-attend to the image representation and autoregressively generate text. While the self-attention layers in the language model remain frozen, the cross-attention layers between the language and vision components are trainable, allowing the model to learn effective multimodal representations.
The VLM’s objective is to predict the next token in a sequence, conditioned on the input image. This is formulated as a cross-entropy loss:
where denotes the image, represents the l-th language token, and represents all tokens preceding . As shown in Figure 3 by combining the components introduced in the previous parts, we can train VLMs either from human annotated data or synthetic image-captions pairs such that the VLM is conditioned on the synthetic image embedding with the cross-entropy loss being between the synthetic captions and the VLM’s predictions.
4 Experiments
4.1 Experimental setup
4.1.1 Datasets
For training we used four datasets: (1) Conceptual Captions v2 (Changpinyo et al., 2021) (CCv2) is a dataset with 12 million image-text pairs that are collected by automatically extracting images and their corresponding “alt-tex” descriptions from various websites. In our experiments we use CCv2 as the main source of human-annotated data whereas for the other datasets we might use synthetically generated images/texts depending on the experiment. (2) WebLI (Chen et al., 2022) has around 350 million image and alt text pairs with high text quality in English. (3) LTIP (Alayrac et al., 2022) has 312 million web collected image and text pairs. (4) GenPair refers to 1 million fully synthetic captions paired with synthetic images generated that we create on-the-fly during the VLM training using our text-to-image generator. The details are described in Section 3.1.
We tested our model on 3 different domains:
Scene description: for this task we used MS-COCO (Chen et al., 2015) and we evaluate the performance of our models on the commonly used test set of MS-COCO under zero-shot and fine-tuning settings. For the finetuning settings we use the training set MS-COCO. We use the Karpathy split (Karpathy and Fei-Fei, 2015) for evaluation. Another dataset we used for this class is Flickr-30k (Plummer et al., 2015): again, we evaluate performances on the Karpathy split. We use CIDEr score (Vedantam et al., 2015) as our metric.
4.1.2 Training details
For both the image generator and the VLM we use a pre-trained and frozen VQ-GAN (see A.1.1 for details on the network). The images are input at a resolution of 256x256, with a patch size of 16, which results in a sequence length of 256 tokens (16x16). Both models use a pre-trained and frozen Chinchilla 400m (Hoffmann et al., 2022) as LLM. The VLM has a perceiver resampler (Alayrac et al., 2022; Jaegle et al., 2021) with 6 layers and 64 latent input queries that cross-attend to the image embeddings. There are cross attention layers between each layer in the LLM and the output of perceiver resampler similar to (Alayrac et al., 2022) ( refer to A.1.4 for details).
We use ViT-B (Dosovitskiy et al., 2020) for the main text-to-image model in the image generator (see Appendix A.1.3 for full details), with a maximum text length of 64 tokens with guidance scale of 4.0, 24 refining steps and a sample choice temperature of 32.5. All the models are trained with AdamW (Loshchilov and Hutter, 2017) with a learning rate of 0.0001, 5000 warmup steps, and the batch size of 512. For finetuning settings (on COCO), we use a learning rate of 0.00001. Section A.1.3 reports additional details. We pre-train the image generator for 500k steps at which point it has a Fréchet inception distance (FID) (Heusel et al., 2017) of 17.1 on MS-COCO test pairs. Our VLMs training experiments all run for 500k steps. Our models are all implemented in JAX and trained on 256 TPUs.
4.2 Results
Method | Real | Synth | #Real Data | #Synth Data | CIDEr-COCO () | CIDEr-Flickr-30 () | VQAV2 Accuracy () | OKVQA Accuracy () |
Baseline | CCv2 | - | 10.1M | - | 22.1 | 12.7 | 29.1 | 32.4 |
Synth2 | CCv2 | LTIP | 10.1M | 330M | 28.7 | 16.7 | 33.5 | 34.5 |
CCv2 | WebLI | 10.1M | 350M | 27.6 | 18.1 | 34.1 | 34.7 | |
CCv2 | WebLI+LTIP | 10.1M | 670M | 33.4 | 20.8 | 35.3 | 36.1 | |
Gold | CCv2+LTIP | - | 340.1M | - | 27.6 | 20.9 | 36.7 | 38.1 |
CCv2+WebLI | - | 360.1M | - | 30.7 | 21.1 | 36.2 | 38.5 | |
CCv2+LTIP+WebLI | - | 690.1M | - | 35.3 | 23.4 | 37.9 | 39.4 |
Real | Synth | #Real Data | #Synth Data | Concentration () | Entropy () | CIDEr-COCO () | Flickr-30() | VQAV2 Acc. () | OKVQA Acc. () |
CCv2 | - | 10.1M | - | - | - | 22.1 | 12.7 | 29.1 | 32.4 |
CCv2 | GenPair | 10.1M | 1M | 57.7% | 3.81 | 25.9 | 17.3 | 31.1 | 34.0 |
CCv2+WebLI | - | 10.1M+1M | - | 69.8% | 3.43 | 24.4 | 14.9 | 30.6 | 33.9 |
CCv2+LTIP | - | 10.1M+1M | - | 83.0% | 2.92 | 23.4 | 13.8 | 30.3 | 32.9 |
4.2.1 Synthetic Images
To first assess the effectiveness of synthetic images compared to the original ones in VLM training, we conduct a study where human-written captions from existing datasets are paired with synthetic images generated by a text-to-image model. We train a VLM on these synthetic image-caption pairs and evaluate its performance against a model trained on the original real image-caption pairs. This comparison allows us to investigate whether synthetic images can effectively substitute for real images in the context of VLM training.
Specifically, we generate synthetic images for the LTIP and WebLI datasets using their annotations. This provides us with (Real Text, Synthetic Image) pairs to train a VLM with (Synth2 in Table 2). We compare this model to the model trained on the human-annotated data of Conceptual Caption v2 (CCv2) only, using no data from LTIP and WebLI (Baseline) and the models trained on the original (Real Text, Real Image) pairs (Gold in Table 2). For consistency, and to reduce model collapse, all Synth2 and Gold models are co-trained with CCv2, the same dataset used by our text-to-image generator and the baseline.
As shown in Table 2, synthetic images significantly improve baseline model performance across the full set of downstream tasks. Importantly, they are effective for VLM training despite the smaller volume of human-annotated images used. Note that “Gold” represents the upper performance bound when using the full original dataset. For LTIP, synthetic images slightly outperform original images, likely due to increased data diversity; while original pairs remain static during training, synthetic images introduce beneficial image augmentation (Azizi et al., 2023).
These results suggest the potential of generative models to enhance VLM training containing complex, multi-object scenes and detailed captions, even with limited real image data.
4.2.2 Synthetic text and image pairs
To further demonstrate the efficacy of using synthetic data for VLM training, we conducted additional experiments where the entire dataset including captions and their corresponding images are generated synthetically by an LLM and the image generator (see Section 2 for the explanation of how these caption where generate and Figure 1D for some examples).
Comparing the first two rows in Table 3 shows that adding even a small fraction (1M) of purely synthetic image-caption data (GenPair) significantly improves performance across all evaluated domains. Interestingly, sampling an additional 1M data points from real datasets like WebLI or LTIP (rows 3 and 4) yields lower improvement.
To investigate this, we assessed the semantic diversity and balance of each caption set. First, we used our language model to generate embeddings for the synthetic GenPair and the real WebLI and LTIP captions. Then, we employed k-means clustering to analyze these embeddings (see Appendix A.3 for details). Figure 4 shows the cluster distribution among all three datasets, with the x-axis representing the cluster index and the y-axis representing data volume from each dataset within a cluster. Notice how LTIP and WebLI have a large concentration of data only within clusters 19, 18, 1, 16 and 0, while most other clusters are under-populated. In contrast, GenPair is distributed more evenly across different clusters (e.g. clusters 1, 4, 6, 9, 12, 14, 15, 16, 17, 18 are all well populated), suggesting superior balance compared to the other two.
To quantify the mentioned semantic concentration within each dataset we use a “Concentration Metric” as the percentage of data belonging to the top-5 most populated clusters. As reported in Table 3, GenPair has the lowest concentration, with only 57.5% of its captions in the top-5 clusters. This indicates lower semantic imbalance in the synthetically generated GenPair and contrasts with the higher imbalance found in the other datasets (69.9% and 83%). Also, the entropy over the full cluster distribution, confirms that GenPair has a more uniform distribution of data across clusters. We postulate that the inherent diversity within GenPair likely contributes to its robust performance on downstream tasks, as models trained on diverse data tend to generalize better.
This evaluation demonstrates Synth2’s capacity to leverage diverse text data, including synthetic data, for enhanced performance. Our proposed analysis and metrics can inform synthetic data creation, further optimizing its impact. Notably, our findings highlight the significant role of synthetic data, generated via image synthesis and LLM captions, in overcoming data limitations and boosting VLM capabilities. This underscores the effectiveness of text-only captions for model enhancement, emphasizing the value of our technique.
4.2.3 Comparison with related work
Furthermore, we investigated the potential of our approach compared to the state-of-the-art model while taking into account the amount of synthetic or real training data and the number of model parameters. We combined all the three datasets namely, GenPair , WebLI and LTIP. Table 4 compares our model’s MS-COCO benchmark performance against related methods: ITIT (Li et al., 2023b), DC (Li et al., 2023c) and SimVLM (Wang et al., 2021). Note that, in this case we provide both zero-shot and finetuning results (see Section A.2 in the Appendix). ITIT employs cycle consistency for VLM improvement, while DC uses Stable Diffusion to generate images based on human-defined heuristics, and SimVLM has one of the state-of-the-art performances on MS-COCO.
Synth2 and ITIT share similar parameter counts and paired data usage, making them directly comparable. As shown in Table 4, Synth2 drastically outperforms ITIT on the CIDEr score. While DC achieves the highest raw performance, it requires significantly more parameters, relies on vast amount of real data, and the filtering procedure is heavily based on human heuristics.
Synth2 strikes a balance between performance and efficiency. By leveraging text-only data, it achieves competitive results using 40 times less paired data than DC. Interestingly, despite performing weaker than SimVLM on finetuning, Synth2 has the highest in zero-shot performance among all models. This comparison highlights the trade-offs between performance, efficiency, and data requirements. Synth2 emerges as a compelling solution when resources are limited, achieving impressive results with reduced data and computational overhead.
4.2.4 Extra analysis
To further analyze the boost in performance afforded by augmentation with fully synthetic data, we characterized the performance of the baseline model (i.e. trained solely on CCv2), and Synth2 which was additionally trained together with GenPair. Figure 5 visualizes the performance trends for both models. Notably, the light blue curve representing the baseline exhibits a gradual increase in performance with increasing training steps, eventually plateauing at 22.1 CIDER score. In contrast, the purple curve reveals Synth2’s steeper performance improvement, achieving comparable performance to the paired only training regime, with roughly 1/3 of the training steps. This highlights the significant data efficiency gain achieved through fully synthetic data augmentation in Synth2. Shaded regions surrounding each curve indicate standard deviation across three random seeds, demonstrating the robustness of the observed performance trends. These findings demonstrate Synth2’s ability to effectively leverage fully synthetic data, resulting in remarkable performance gains.
Our study also investigated the computational efficiency of Synth2. We compared the case where captions are used to generate image embedding versus when captions are rendered into actual images. As shown in Table 5, Synth2 trained using image embedding consistently demonstrates faster performance, running at 2.08 training steps per second, whereas Synth2 trained from pixel runs at a slower pace of 1.66 steps/second. Critically, the performance on image captioning and visual question answering are not affected. This efficiency advantage stems from Synth2’s utilization of text-only data and embedding-based image generation using parallel decoding, which reduces the computational overhead associated with processing pixel-based images as done in the Synth2 model trained from pixels without affecting the performance. For further studies, e.g. comparing the performance of different image encoder backbones refer to Appendix, Table 9.
Model | All Params | Real Data | Synth Data | CIDEr (FT) | CIDEr (zs) |
ITIT | 11.2B | 3M | 110M | 103.5 | 32.1 |
DC (BLIP) | 1.7B | 5.5B | 400M | 133.1 | - |
SimVLM | 1.4B | 1.1B | 365M | 143.3 | 32.2 |
Synth2 | 632M | 10.1M | 711M | 131.3 | 35.4 |
Model | Step/Sec | MSCOCO CIDEr(zs) | VAQV2 Accuracy (ft) |
Synth2- Embeddings | 2.08 | 25.9 | 50.1 |
Synth2- Pixels | 1.66 | 26.0 | 49.9 |
5 Limitations
The two main limitations of the current work stem from the restricted quantity of fully synthetic data used and the limited exploration of text data sources. Firstly, whilst we were able to show a substantial gain when training was augmented with fully synthetic data, our experiments were limited to a relatively low quantity of data (i.e. 1M image-caption pairs). In the future, it will be important to examine whether using orders of magnitude more fully synthetic data (e.g. 700M) will result in performance gains that surpass that of using solely paired data. Also, the current work primarily explores a restricted set of text-only data sources. Further research is necessary to investigate the potential of diverse text sources, including investigating context-specific generation (e.g., medical data), which could be an exciting avenue for future exploration. By acknowledging and addressing these limitations, future research can leverage the strengths of the Synth2 approach and enhance its generalizability, robustness, and efficiency for a broader range of visual-language modeling tasks.
6 Conclusions
Our study presents a novel approach for generating synthetic image-text pairs to enhance the training of visual language models. By harnessing the capabilities of large language models and text-to-image generation, we effectively address the limitations of manual image labeling such as scalability, customization and costs. In particular, we show that the visual language model trained on our synthetic and human annotated datasets exhibits a marked improvement in image captioning tasks compared to a baseline trained exclusively on human-annotated data. Overall, this research introduces a novel technique that has the potential to transform visual language model training. Our findings highlight the value of synthetic data generation, paving the way for advancements in numerous fields where visual language understanding is crucial.
7 Broader Impact
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here. In particular biases in generative models can have a drastic effect on the synthetic data. Similarly it is important to consider the challenges of privacy when using generative models.
Acknowledgements
We would like to thank Sander Dieleman, Ali Razavi, and Benigno Uria for their insights on VQGAN. We would also like to thank Aida Nematzadeh, Pinelopi Papalampidi, Han Zhang, Mateusz Malinowski, Valentin De Bortoli, Sahra Ghalebikesabi, Emanuele Bugliarello, Chris Knutsen, and Murray Shanahan for their in depth comments and support throughout the project.
Appendix A Appendix.
A.1 Extra Implementation details
A.1.1 VQ-GAN details
Config param | Value |
Perceptual loss weight | 0.05 |
Adversarial loss weight | 0.01 |
Codebook size | 8192 |
Optimizer | Adam (Kingma and Ba, 2014) |
Discriminator learning rate | 1e-4 |
Generator learning rate | 1e-4 |
Optimizer momentum | =0.9 =0.99 |
Batch Size | 256 |
Learning rate schedule | Cosine |
Decay (Loshchilov and Hutter, 2017) Warmup steps | 10000 |
Training steps | 500000 |
VQGAN Architecture: Our VQGAN architecture is similar to the previous work (Esser et al., 2021). It consists of several residual blocks, downsample (encoder) and upsample (decoder) blocks. The main difference is that we remove the non-local block to make the encoder and decoder fully convolutional to support different image sizes. In the base VQGAN model, we apply 2 residual blocks in each resolution and the base channel dimension is 128. For the finetuned decoder, we apply 4 residual blocks in each resolution and we also make the base channel dimension to be 256.
A.1.2 Optimization details for Synth2
Config param | Value |
Optimizer | AdamW (Loshchilov and Hutter, 2017) |
Learning rate | 1e-4 |
Warmup steps | 5e3 |
Weight Decay | 1e-4 |
Gradient Clip | 1.0 |
A.1.3 Text to image generator details
Config param | Value |
Num Layers | 12 |
Num Heads | 12 |
Embedding hidden dim | 768 |
MLP hidden dim | 3072 |
Dropout | 0.1 |
Codebook size | 8192 |
Max Sample Length | 64 |
Guidance Scale | 4 |
Sample Temperature | 32.5 |
Activation | GeLU |
A.1.4 Perceiver details
Config param | Value |
Num Layers | 6 |
Num Heads | 16 |
Embedding hidden dim | 1024 |
Activation | GeLU |
A.2 Finetuning details
Following the pre-training stage, our model undergoes fine-tuning for various downstream tasks. The fine-tuning process employs the AdamW optimizer with similar Beta values as in pre-training. The learning rate was 1e-5.
To enhance generalization capabilities during fine-tuning, we utilize regularization methods such as Dropout (set to 0.1).
In accordance with established practices, we use the development split to determine optimal fine-tuning settings. The performance results are then reported based on the test split.
A.3 Clustering details
To determine the optimal number of clusters for our analysis, we employed the Elbow Method in conjunction with the Within-Cluster Sum of Squares (WCSS) metric. We iteratively applied the K-means clustering algorithm while varying the number of clusters, calculating the WCSS for each iteration. A clear ‘elbow’ point was observed in the WCSS plot, indicating a substantial decrease in variance as the number of clusters increased up to 20. Beyond this point, further increases in the number of clusters yielded diminishing returns in terms of WCSS reduction. Based on this analysis, we determined that 20 clusters provided a suitable balance between parsimony and capturing the underlying structure within our dataset.
To analyze the diversity of cluster compositions, we employed a co-clustering approach with the K-means algorithm on the concatenated GenPair, WebLI, and LTIP datasets. For each of the resulting clusters, we calculated the normalized cluster sizes for each individual dataset. This allowed us to visualize the distribution among the datasets within each cluster, as illustrated in Figure 4.
To calculate the concentration we performed a cumulative sum over the top-5 clusters to determine what percentage of the data points was present there. The idea is that, the higher the percentage in the top 5 clusters, the less uniform is the distribution across all clusters, potentially indicating a lower level of diversity. Here in Table 8 we also report top-3 and the entropy over the full 20 clusters. Both measures confirm a more uniform distribution of data across clusters for GenPair versus the WebLi and LTIP. Especially the entropy complement the other measures as it takes into account all clusters, even if they are smaller.
Real | Synth | #Real Data | #Synth Data | Concentration Top-3 () | Concentration Top-5 () | Entropy () | CIDEr-COCO () | Flickr-30() | VQAV2 Acc. () | OKVQA Acc. () |
CCv2 | - | 10.1M | - | - | - | - | 22.1 | 12.7 | 29.1 | 32.4 |
CCv2 | GenPair | 10.1M | 1M | 40.5% | 57.7% | 3.81 | 25.9 | 17.3 | 31.1 | 34.0 |
CCv2+WebLI | - | 10.1M+1M | - | 58.2% | 69.8% | 3.43 | 24.4 | 14.9 | 30.6 | 33.9 |
CCv2+LTIP | - | 10.1M+1M | - | 76.5% | 83.0% | 2.92 | 23.4 | 13.8 | 30.3 | 32.9 |
A.4 Backbone Comparison
We conducted an ablation study to compare the performance of different visual backbones. For this comparison, we kept the total amount of FLOPs (both pre-training and training) and (self)supervised training data constant across models. We compared our VQ-GAN-based backbone against a contrastively pre-trained NFNet-F6 model (Brock et al., 2021) using the same loss function as CLIP (Radford et al., 2021). As shown in Table 9, the primary factor influencing performance appears to be the number of parameters, as our model with a deeper VQ-Adapter (VIT-L+) achieved results comparable to the contrastively trained model. In our study, we opted for the Synth2VIT-B backbone due to its compatibility with our hardware resources. This choice allowed us to avoid model sharding, thereby enabling faster experimental iterations.
Model | real data | #params | MSCOCO CIDEr | VAQV2 Accuracy |
Synth2 VIT-B adapt. | 690.1 | 632M | 35.3 | 37.9 |
Synth2 VIT-L+ adapt. | 690.1 | 890M | 40.9 | 42.8 |
NFNet-F6 (CLIP) | 690.1 | 920M | 41.3 | 43.4 |
A.5 Image Generator Qualitatives